OpenStudy (anonymous):

Find a power series representation for the function and determine the interval of convergence. See first post for the problem, second post will be work i have so far.

OpenStudy (anonymous):

\[f(x)=\frac{ 5 }{ 2-4x ^{2} }\]

OpenStudy (anonymous):

\[\large \frac{ 5 }{ 2 }\frac{ 1 }{ 1-2x ^{2} }=\sum_{n=0}^{\infty} \frac{ 5 }{2 } (2x ^{2})^{n}\] This is where i'm not sure if I am doing this right. All the examples I could find did not have the a different constant to multiply back in.

terenzreignz (terenzreignz):

Don't worry, I don't really see any errors...

terenzreignz (terenzreignz):

I suppose you'll have no problem with finding the interval of convergence?

OpenStudy (anonymous):

well i had found it and realized i forgot the 5/2 do i just multiply the 5/2 to the 2x^2 and solve?

terenzreignz (terenzreignz):

No, the constant 5/2 has nothing to do with the interval of convergence. It's just a constant factor after all, so it won't affect convergence one bit. Now, as for the interval of convergence, have you found it?

OpenStudy (anonymous):

well if that does not interfere then I have: \[\large \frac{ -1 }{ \sqrt{2} }\le x \le \frac{ 1 }{\sqrt{2} }\]

terenzreignz (terenzreignz):

nu-uh :) Drop the equality, you know full well that if x is equal to any of those two endpoints, the resulting series diverges :P

OpenStudy (anonymous):

sorry the paper had it without the equal, was using those helping someone else: \[\frac{ -1 }{ \sqrt{2} }<x <\frac{ 1 }{ \sqrt{2} }\] so that would be right?

terenzreignz (terenzreignz):

Yup. It converges whenever x is within that interval. Good job ^_^

OpenStudy (anonymous):

ok thank you